OpenStudy (anonymous):

Will give fan and medal. How long would the rocket take to fall back to Earth? A model rocket is launched from the ground with an initial velocity of 160 ft/sec. How long will it take the rocket to reach its maximum height? Show all work in the space provided. I worked up the equation myself Height = f(t) = 160t – 16t² f(1) = 144 ft. Velocity = f`(t) = 160 – 32t f`(1) = 128 ft/s. f`(t) = 0 at the maximum height 0 = 160 – 32t 32t = 160 t = 5 5 seconds f(t) = 160t – 16t² f(5) = 160(5) – 16(5)² f(5) = 800

OpenStudy (anonymous):

So with this equation done, how long would it take for the rocket to fall back to earth from the time it was launched if it's Parachute failed to deploy?

OpenStudy (anonymous):

@phi can you help me?

OpenStudy (welshfella):

you can use one of the equations of motion for constant acceleration. In this case the acceleration is due to gravity and = 32 ft /s^2 v = u + at at maximum height v = final velocity = 0 , u = 160 and a = -32 ( as the acceleration acts in opposite direction to motion) 0 = 160 - 32t t = 160/32 secnds

OpenStudy (welshfella):

t = 5 seconds as you said

OpenStudy (anonymous):

@welshfella can I ask another question?